A charge of -5.16 μC is traveling at a speed of 5.36 × 106 m/s in a region of space where there is a magnetic field. The angle between the velocity of the charge and the field is 53.3o. A force of magnitude8.99 × 10-3 N acts on the charge. What is the magnitude of the magnetic field?
Answer #1

Given:

q=5.16\times10^{-6}\ \text C\\ v=5.36\times10^6\ \text{m/s}\\ \theta=53.3^0\\ F=8.99\times10^{-3}\ \text N

Using the Lorentz force law, the magnetic field is,

B=\frac{F}{qv\sin\theta}\\ =\frac{8.99\times10^{-3}}{5.16\times10^{-6}\times5.36\times10^6\times\sin(53.3)}\ \text T\\ =4.05\times10^{-4}\ \text T

Note:

Here the polarity of the charge is neglected because we need only magnitude of the field.